OpenStudy (hesan):

If (A+iB)=(2+i)/(2-3i), then A^2+B^2=?

OpenStudy (anonymous):

pout the values

OpenStudy (hesan):

values of what?

OpenStudy (hesan):

what is the answer?

OpenStudy (zepp):

We have this, factored by using complex numbers \[\large a^2 + b^2 = (a + bi)(a - bi)\]

OpenStudy (zepp):

You have \((a+bi)\) and \((a-bi)\) is its conjugate.

OpenStudy (hesan):

ok, the answer i m getting is 5/13. is this correct?

OpenStudy (zepp):

\[(\frac{2+i}{2-3i})(\frac{2-i}{2+3i})\] I believe it's something like this, someone can confirm? I don't work often with complex numbers :[

ganeshie8 (ganeshie8):

i got the 5/13 @zepp numerator also 2+3i... rationalizing denom right ?

OpenStudy (hesan):

ya right, the answer is 5/13 then..

OpenStudy (zepp):

If I'm correct, it would become \[\large \frac{2^2-i^2}{2^2-(3i )^2}=\frac{4-(-1)}{4-9(-1)}=\frac{4+1}{4+9}=\frac{5}{13}\] That's my work

OpenStudy (hesan):

Great! Thnx!

OpenStudy (zepp):

You are welcoe :)

OpenStudy (zepp):

welcome*

OpenStudy (shubhamsrg):

no need to rationalize

ganeshie8 (ganeshie8):

yeah... i see that lol i was thinking of rationalizing and comparing i terms & real ones...

OpenStudy (shubhamsrg):

its asking you to take modulus of both sides : use them : 1)|a + ib| = sqrt(a^2 + b^2) 2)|p/q| = |p|/|q| where p and q are complex nos..

mathslover (mathslover):

rather than these answers i will prefer u to go to these links also : 1) for practice : http://openstudy.com/study#/updates/4fe5a252e4b06e92b873c296 2) for learning : http://openstudy.com/study#/updates/4fe196d3e4b06e92b870a14a best of luck @Hesan